OpenStudy (anonymous):

add: v/v^2-5v-6 - 1/v+1.. i know that if you mulitply the second fraction by v-6 it will give you the same denomitator i think.. but i need help please

OpenStudy (anonymous):

Is your question,\[\frac{v}{v^2}-5v-6-\frac{1}{v}+1\]?

OpenStudy (anonymous):

I'm going to type out for this question. If my interpretation is wrong, you're best to re-post. \[\frac{v}{v^2}-5v-6-\frac{1}{v}+1\]\[=\frac{1}{v}-5v-6-\frac{1}{v}+1\]\[=-5v-5\](since the 1/v's cancel)\[=-5(v+1)\]

OpenStudy (anonymous):

how do you get the fraction

OpenStudy (anonymous):

on top of the first fraction is v on the bottom is v^2-5v-6 + on top of the second is 1 on bottoom is v+1

OpenStudy (anonymous):

Okay, you need to use brackets to group things.

OpenStudy (anonymous):

i do not know how to write equations on here... i do not understand i type it in the best i can

OpenStudy (anonymous):

[(v \ v ^{2}-5v-6) (1 \ v+1)]

OpenStudy (anonymous):

in the middle it is subtract.. between the two fractions

OpenStudy (anonymous):

nvm it is add. so ill re do it

OpenStudy (anonymous):

\[(v \ v^2-5v-6)+(1 \ v+1)\]

OpenStudy (anonymous):

lol, you're really struggling with this editor

OpenStudy (anonymous):

between the v and v^2 the V^2 part should be on bottom and the v+1 should be on the bottom

OpenStudy (anonymous):

okay...wait...

OpenStudy (anonymous):

i get it

OpenStudy (anonymous):

\[\frac{v}{v^2-5v-6}+\frac{1}{v+1}\]is this it?

OpenStudy (anonymous):

abarnzy, still there? Is that expression right before I go on?

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

You can first factor the denominator of the first fraction as\[v^2-5v-6=(v-6)(v+1)\]to get

OpenStudy (anonymous):

\[\frac{v}{(v-6)(v+1)}+\frac{1}{v+1}=\frac{1}{v+1}\left[ \frac{v}{v-6}+1 \right]\]\[=\frac{1}{v+1}\left[ \frac{v+(v-6)}{v-6} \right]=\frac{2v-6}{(v+1)(v-6)}=\frac{2(v-3)}{(v+1)(v-6)}\]